# Machine Space I: Weak exponentials and quantification over compact spaces

### v3

@@ Abstract.
Topology may be interpreted as the study of verifiability, where opens correspond to semi-decidable properties. In this paper we make a distinction between verifiable properties themselves and processes which carry out the verification procedure. The former are simply opens, while we call the latter machines. Given a frame presentation $\mathcal{O}X=\langle G\mid R\rangle$ we construct a space of machines $\Sigma^{\Sigma^{G}}$ whose points are given by formal combinations of basic machines corresponding to generators in $G$. This comes equipped with an ‘evaluation’ map making it a weak exponential for $\Sigma^{X}$.
@@ 1. Introduction
Topology may be viewed as the logic governing verifiability. A property $P$ is said to be verifiable if whenever an element satisfies $P$, this can be established by finite means. Notably absent are any constraints on elements which do not satisfy $P$ — that is, you need not be able to verify this fact. This perspective is due to Smyth in [22, 23] and has been expanded on subsequently in [28] and various other papers discussed below. In this view open sets are interpreted as collections of points satisfying verifiable properties — with intersection and union corresponding to logical conjunction and disjunction, respectively. This viewpoint motivates the axioms of a topological space and continuous functions. It is possible to then derive interpretations of various other topological concepts in terms of verifiability. In this paper we make use of the pointfree approach which emphasises the verifiable properties $P$ over the collections of points which satisfy them. The analogue of topological spaces from this perspective are called locales, though we will often simply call these ‘spaces’ when speaking intuitively.
@@
We might be tempted to identify machines with opens, which can in turn be seen to correspond to continuous maps from $X$ into the Sierpiński space $\Sigma$. Since we are thinking of these machines as ‘real things’ we can interact with, we can expect there to be natural verifiable properties concerning the machines themselves (for example, does the machine halt on some given point) yielding a space of machines for each space $X$. However, if $X$ is not locally compact, the exponential object $\Sigma^{X}$, which we would think of as the space of opens, does not exist and so we are forced to distinguish between machines and opens — unlike functions from $X$ to $\Sigma$, machines are not extensional. Moreover, since spaces of machines should always exist, machines are better than opens for some purposes.
In order the formalise these ideas, we describe an explicit construction of a space of machines. We replace the (nonexistent) space of opens $\Sigma^{X}$ with a certain weak exponential. Our starting point is to fix a presentation for $X$ with generators $G$. (From the classical perspective these can be viewed as a set of subbasic opens $G$ for the space $X$.) Since $G$ is a set, the space $\Sigma^{G}$ exists. The space $X$ embeds naturally into $\Sigma^{G}$ and so the opens of $X$ are restrictions of opens in $\Sigma^{G}$. We will view distinct opens of $\Sigma^{G}$ which restrict to the same open $U$ in $X$ as distinct machines which accept (i.e. halt on) precisely the elements of $U$. A crucial point is that the space $\Sigma^{G}$ is locally compact and we can take $\Sigma^{\Sigma^{G}}$ to be the space of machines of $X$. This space may be thought of as a reasonably canonical weak exponential associated to $X$ with base $\Sigma$. More concretely, the points of $\Sigma^{\Sigma^{G}}$ can be thought of as formal joins of formal meets of the generators, which are represented by certain programs that run the ‘basic machines’ from $G$ in parallel.
We then relate machine space to $\Sigma^{X}$, when the latter exists. Under the interpretation so far described an open merely represents a verifiable property in an abstract sense, whereas a machine is some process that concretely semi-decides memberships of elements. Given an open and a point in a general space $X$ there is not an obvious way to verify that the point lies in the open. This helps explain the fact that for general spaces the collection of opens equipped with the Scott topology does not have a continuous evaluation map. We could however expect the evaluation map to be continuous if there were some way to associate a machine to each open. Indeed, we show that when $X$ is locally compact the canonical quotient map of machine space onto $\Sigma^{X}$ always has a section, allowing a continuous assignment of opens to machines which represent them.
@@
If $K$ is a compact, locally compact space, then we can check if a verifiable property holds on all of $K$ and so we should be able to verify whether the corresponding open in $\Sigma^{K}$ is equal to the largest element of $\Sigma^{K}$, namely $K$ itself. In other words, $K$ is compact if and only if the singleton $\{K\}$ is open in $\Sigma^{K}$. See [6]. Of course, this viewpoint is inapplicable outside of the locally compact case, since the exponential $\Sigma^{K}$ will not exist.
As might be expected, this can be extended to the general setting by replacing $\Sigma^{K}$ with machine space. This perspective on the universal quantification is essentially due to Escardó [5]. Instead of asking if an open of $K$ equals $K$, we ask if an open in $\Sigma^{G}$ contains $K$ — that is, if the machine accepts all points of $K$. By the Hofmann–Mislove theorem [11], there is an associated open corresponding to all of the machines which cover $K$, which plays the same role as the singleton $\{K\}$ in the previous approach.
Our contribution is obtaining this open via an explicit algorithm for universally quantifying over a compact space. This is a very general and purely topological version of Escardó’s algorithm [5, 7] for universal quantification over Cantor space. (See also [8] which discusses universal quantification over data types in the setting of higher type computation.) Unlike Escardó, we do not work within a particular formal semantics, but remain agnostic towards the particular ‘model’ of the topological structures.
@@ 2. Background
2.1. The verifiable interpretation of topology
We begin by discussing the interpretation of topology in terms of verifiable properties. For more details, see [23, 28].
Let $X$ be some collection of points and let us consider what properties the set $\mathcal{O}X$ of verifiable propositions about $X$ should have. Given two propositions $U$ and $V$ we can verify their conjunction $U\wedge V$ by simply verifying $U$ and $V$ in turn. Furthermore, the constantly true proposition is trivially verifiable and so $\mathcal{O}X$ is closed under finite conjunctions. Note that this argument does not extend to infinite conjunctions, since checking infinitely many verifiable propositions in turn is not achievable via finite means. However, verifiable properties are closed under disjunctions of arbitrary cardinality: to verify that a disjunction holds at some point, we need only verify that one disjunct holds, which does not require any infinite processes. Note in particular that the constantly false position is verifiable, since we need only verify propositions when they hold.
It is instructive to consider why such reasoning cannot be applied to prove that every physically realisable function $f\colon X\to Y$ is open — that is, that the image of every open $V$ in $X$ is open in $Y$. How would we verify that that $y\in f(V)$? We would need to semi-decide that $y=f(x)$ for some $x\in V$. But equality is not verifiable in general and so we are not always able do this. If we could semi-decide equality, then each singleton $\{y\}$ would necessarily be open, which would make $Y$ a discrete space, and indeed, all functions into discrete spaces are open.
Example 2.1.
Consider the case of the real numbers $\mathbb{R}$. We might imagine having devices that can measure a quantity $x$, each to some fixed precision giving us a rational number. Moreover, let us assume that for any precision there exists a device that is at least that precise.
@@ 2.2. Pointfree topology
The logic of verifiability can also be studied purely through the verifiable propositions themselves, without any regard to the points which may or may not satisfy these propositions. This is perspective of pointfree topology, the study of topology through the lattice of open sets. In this section we introduce some of the basic notation and concepts which will be used throughout the rest of the paper. See [17, 28] for further details.
@@
$u\wedge\bigvee_{i\in I}v_{i}=\bigvee_{i\in I}u\wedge v_{i}.$
@@
Of course, the open sets of any topological space form a frame under intersection and union, though there are also non-spatial examples.
@@
Since frames are algebraic structures, they can be presented by generators and relations. In particular, free frames exist. The free frame on $G$ can be described explicitly as the frame of downsets on the meet-semilattice of finite subsets of $G$ ordered by reverse inclusion. The finite subsets of $G$ are interpreted as a formal meets of generators, while the downsets are viewed as formal joins of these. Note that every formal expression involving frame operators on the generators can be brought into the form of a join of finite meets by distributivity.
@@
A presentation for a frame consists of a set of generators $G$ and a set of relations $R$ consisting of formal equalities (or inequalities) between formal combinations of generators — for example, $g_{1}\wedge g_{2}=\bigvee_{i=3}^{\infty}g_{i}\wedge g_{2}$. Explicitly, $R$ can be defined to be a subset of $F(G)\times F(G)$ where $F(G)$ is the free frame on the set $G$.
@@
[…]
Note that if $G$ is a set then $\Sigma^{G}$ is the locale corresponding to the free frame on $G$ and so a presentation of $X$ can be understood as an equaliser of locales $X\hookrightarrow\Sigma^{G}\rightrightarrows\Sigma^{R}$. (Recall that locales are dual to frames and so a coequaliser of frames becomes and equaliser of locales.)
@@
As mentioned before, compactness has a very nice interpretation in terms of universal quantification of verifiable properties. We will see some instances of this in later sections.
@@ 2.3. The Scott topology and exponentials
A subset $V\subseteq\mathcal{O}X$ is Scott-open if it is upward closed and if whenever a join $\bigvee D\in V$ for a directed set $D$, then some $d\in D$ lies in $V$. Recall that a set $D$ is directed if every finite subset of $D$ has an upper bound in $D$. ∎
If $X$ is locally compact (see Definition 3.16 for a formal definition), then $X$ is exponentiable (see [12]). In particular, the exponential locale $\Sigma^{X}$ exists and coincides with the Scott topology on $\mathcal{O}X\cong\mathrm{Hom}(X,\Sigma)$. Furthermore, $\Sigma^{X}$ is itself locally compact. However, if $X$ is not locally compact the evaluation map $(f,x)\mapsto f(x)$ is not continuous with respect to the Scott topology and the exponential $\Sigma^{X}$ does not exist.
The Scott topology can actually be defined more generally on dcpos, which are often used to model data types in programming languages.
@@
A poset which admits joins of all directed subsets is called a dcpo. We will write $\operatorname*{\bigvee{}^{\!\uparrow}}S$ for the join of a directed set $S$. A morphism of dcpos or a Scott-continuous function is a monotone map which preserves directed suprema. ∎
The order on a dcpo can be thought of as a definability order: the least element (if it exists) corresponds to a purely divergent computation, while maximal elements are completely defined. For more on this topic see [1, 2, 10].
@@ 3. Machine space
In this section we make certain aspects of the verifiable interpretation of topology more precise, specifically the relationship between opens and the processes that semi-decide some fact. Our approach makes use of presentations of frames and inspires a number of purely topological results — many with interesting interpretations from the point of view of verifiability.
Given a space $X$ with $\mathcal{O}X=\langle G\mid R\rangle$ we interpret each generator $g\in G$ as a basic black-box machine which takes as input the points of $X$ and either halts after some finite amount of time, or runs forever. More complicated machines can be constructed from these basic machines. For instance, we may construct a composite machine by taking a finite collection of machines and insist that each machine halts on the given point before the composite machine is considered to have halted. We can also construct a further machine by running any number of these composite machines in parallel and halting on a point if at least one branch halts. Topologically this corresponds to taking formal joins of finite formal meets of generators.

### v4

@@ Abstract.
Topology may be interpreted as the study of verifiability, where opens correspond to semi-decidable properties. In this paper we make a distinction between verifiable properties themselves and processes which carry out the verification procedure. The former are simply opens, while we call the latter machines. Given a frame presentation $\mathcal{O}X=\langle G\mid R\rangle$ we construct a space of machines $\Sigma^{\Sigma^{G}}$ whose points are given by formal combinations of basic machines corresponding to generators in $G$. This comes equipped with an ‘evaluation’ map making it a weak exponential with base $\Sigma$ and exponent $X$.
@@ 1. Introduction
The interpretation of topology in terms of verifiable properties provides insight into many aspects of the field. However, if this perspective is taken completely seriously it prompts some tricky questions concerning the failure of exponentials of topological spaces. In this paper, we will address these questions by introducing the notion of machine space and demonstrate its utility by extending Escardó’s algorithm for universal quantification over compact spaces (see [9]) from data types to general spaces.
A property $P$ is said to be verifiable if whenever an element satisfies $P$, this can be established by finite means. Notably absent are any constraints on elements which do not satisfy $P$ — that is, you need not be able to verify this fact.
Let $X$ be some collection of points and let us consider what properties the set $\mathcal{O}X$ of verifiable propositions about $X$ should have. Given two propositions $U$ and $V$ we can verify their conjunction $U\wedge V$ by simply verifying $U$ and $V$ in turn. Furthermore, the constantly true proposition is trivially verifiable and so $\mathcal{O}X$ is closed under finite conjunctions. Note that this argument does not extend to infinite conjunctions, since checking infinitely many verifiable propositions in turn is not achievable via finite means. However, verifiable properties are closed under disjunctions of arbitrary cardinality: to verify that a disjunction holds at some point, we need only verify that one disjunct holds, which does not require any infinite processes. Note in particular that the constantly false proposition is vacuously verifiable, since we need only verify propositions when they hold.
Example 1.1.
Consider the case of the real numbers $\mathbb{R}$. We might imagine having devices that can measure a quantity $x$, each having some fixed precision $\varepsilon>0$ and giving rational outputs. More precisely, for a real number $x$, such a device will (nondeterministically) give some $q\in\mathbb{Q}$ such that $\lvert x-q\rvert<\varepsilon$. Moreover, let us assume that for any precision there exists a device that is at least that precise.
This interpretation of topology is due to Smyth in [22, 23] and has been expanded on subsequently in [27] and various other papers discussed below.
Given this perspective, it is possible to derive interpretations of various other topological concepts in terms of verifiability. For instance, it is well known that a compact set $K$ corresponds to a set that can be universally quantified over — that is, given a verifiable property $P$ it is possible to verify that each member of $K$ satisfies $P$ (Note that $K$ may in fact be infinite!)
In this paper we make use of the pointfree approach which emphasises the verifiable properties $P$ over the collections of points which satisfy them. The analogue of topological spaces from this perspective are called locales, though we will often simply call them ‘spaces’.
Philosophical problems
The verifiability interpretation raises the following problem of a philosophical nature: do the verifiable properties on a space themselves have verifiable properties so that they can be arranged into a topological space? Naively the answer would appear to be ‘yes’. For instance, let $x\in X$ be a point in a space $(X,\mathcal{O}X)$. It appears that we should be able to verify whether a verifiable property in $\mathcal{O}X$ contains $x$ or not — as for any verifiable property $U$, by definition we can verify if an element $x$ satisfies it. While this is a compelling argument, the mathematics tells a different story. What we are asking for is equivalent to asking for a certain exponential object to exist in the category of topological spaces, but it is well known that these exponentials do not always exist. How can we reconcile this failure with our philosophical intuition?
A seemingly unrelated problem, which was actually the original motivation behind the paper, concerns compactness: if compact spaces allow for universal quantification of open predicates, is there a uniform procedure for universal quantification that works for all compact spaces? Or do different spaces require ad-hoc procedures?
The work of Escardó indicates that it is likely the former. In [9, 10] he provides such a uniform algorithm for decidable properties in the context of programming language semantics. While this is compelling, it is not the complete solution to our problem. For one, Escardó is mainly interested in decidable properties, instead of the more general verifiable properties. Moreover, these papers concern Kleene–Kreisel spaces instead of general topological spaces or locales.
Both of these problems are concerned with verifiable properties themselves and in what follows we provide a treatment of them that resolves both problems.
Mathematical solutions
@@
We might be tempted to identify machines with opens, which can in turn be seen to correspond to continuous maps from $X$ into the Sierpiński space $\Sigma$. Since we are thinking of these machines as ‘real things’ we can interact with, we can expect there to be natural verifiable properties concerning the machines themselves (for example, does the machine halt on some given point?) yielding a space of machines for each space $X$. However, if $X$ is not locally compact, the exponential object $\Sigma^{X}$, which we would think of as the space of opens, does not exist and so we are forced to distinguish between machines and opens — unlike functions from $X$ to $\Sigma$, machines are not extensional. Moreover, since spaces of machines always exist, machines are better than opens for some purposes.
In order to formalise these ideas, we describe an explicit construction of a space of machines. We replace the (potentially nonexistent) space of opens $\Sigma^{X}$ with a certain weak exponential. Our starting point is to fix a presentation for $X$ with generators $G$. (From the classical perspective these can be viewed as a set of subbasic opens $G$ for the space $X$.) Since $G$ is a set, the space $\Sigma^{G}$ exists. The space $X$ embeds canonically into $\Sigma^{G}$ and so the opens of $X$ are restrictions of opens in $\Sigma^{G}$. We will view distinct opens of $\Sigma^{G}$ which restrict to the same open $U$ in $X$ as distinct machines which accept (i.e. halt on) precisely the elements of $U$. A crucial point is that the space $\Sigma^{G}$ is locally compact and we can take $\Sigma^{\Sigma^{G}}$ to be the space of machines of $X$. This space may be thought of as a reasonably canonical weak exponential associated to $X$ with base $\Sigma$. More concretely, the points of $\Sigma^{\Sigma^{G}}$ can be thought of as formal joins of formal meets of the generators, which are represented by certain programs that run the ‘basic machines’ from $G$ in parallel.
We then relate machine space to $\Sigma^{X}$, when the latter exists. In particular, there is a canonical quotient map $\Sigma^{\Sigma^{G}}\twoheadrightarrow\Sigma^{X}$ sending each machine to its associated open. Under the interpretation so far described an open merely represents a verifiable property in an abstract sense, whereas a machine is some process that concretely semi-decides memberships of elements. Given an open and a point in a general space $X$ there is not an obvious way to verify that the point lies in the open, as the open alone does not give any such procedure. This helps explain the fact that for general spaces the collection of opens equipped with the Scott topology does not have a continuous evaluation map. We could however expect the evaluation map to be continuous if there were some way to associate a machine to each open. Indeed, we show that when $X$ is locally compact the canonical quotient map of machine space onto $\Sigma^{X}$ always has a section, allowing a continuous assignment of opens to machines which represent them.
@@
If $K$ is a compact, locally compact space, then we can check if a verifiable property holds on all of $K$ and so we should be able to verify whether the corresponding open in $\Sigma^{K}$ is equal to the largest element of $\Sigma^{K}$, namely $K$ itself. In other words, $K$ is compact if and only if the singleton $\{K\}$ is open in $\Sigma^{K}$ (see [8]). Of course, this viewpoint is inapplicable outside of the locally compact case, since the exponential $\Sigma^{K}$ will not exist [13].
As might be expected, this can be extended to the general setting of compact (but not necessarily locally compact) locales by replacing $\Sigma^{K}$ with machine space. This perspective on the universal quantification is essentially due to Escardó [7]. Instead of asking if an open of $K$ equals $K$, we ask if an open in $\Sigma^{G}$ contains $K$ — that is, if the machine accepts all points of $K$. By the Hofmann–Mislove theorem [12], there is an associated open corresponding to all of the machines which cover $K$, which plays the same role as the singleton $\{K\}$ in the previous approach.
Our contribution is obtaining this open via an explicit algorithm for universally quantifying over a compact space. This is a very general and purely topological/localic version of Escardó’s algorithm [7, 9] for universal quantification over Cantor space. (See also [10] which discusses universal quantification over data types in the setting of higher type computation.) Unlike Escardó, we do not work within a particular formal semantics, but remain agnostic towards the particular ‘model’ of the topological structures.
@@ 2.1. Pointfree topology
The logic of verifiability can also be studied purely through the verifiable propositions themselves, without any regard to the points which may or may not satisfy these propositions. This is the perspective of pointfree topology, the study of topology through the lattice of open sets. In this section we introduce some of the basic notation and concepts which will be used throughout the rest of the paper. See [19, 27] for further details.
@@
$u\wedge\bigvee_{i\in I}v_{i}=\bigvee_{i\in I}(u\wedge v_{i}).$
@@
Of course, the open sets of any topological space form a frame under intersection and union. Examples of this form are called spatial. There are also non-spatial examples (such as complete atomless Boolean algebras).
@@
Frames are well-behaved (infinitary) algebraic structures and can be presented by generators and relations (see [14]). In particular, free frames exist. The free frame on $G$ can be described explicitly as the frame of downsets on the meet-semilattice of finite subsets of $G$ ordered by reverse inclusion. The finite subsets of $G$ are interpreted as formal meets of generators, while the downsets are viewed as formal joins of these. Note that every formal expression involving frame operators on the generators can be brought into the form of a join of finite meets by distributivity.
@@
A presentation for a frame consists of a set of generators $G$ and a set of relations $R$ consisting of formal equalities (or inequalities) between formal combinations of generators — for example, $g_{1}\wedge g_{2}=\bigvee_{i=3}^{\infty}g_{i}\wedge g_{2}$. Explicitly, $R$ can be defined to be an (arbitrary) subset of $F(G)\times F(G)$ where $F(G)$ is the free frame on the set $G$.
@@
A general presentation $\langle G\mid R\rangle$ can be obtained from a coequaliser $F(R)\rightrightarrows F(G)\twoheadrightarrow\langle G\mid R\rangle$ where the maps $F(R)\rightrightarrows F(G)$ are obtained by the universal property of $F(R)$ from the natural maps $R\rightrightarrows F(G)$.
One particularly important locale is $\Sigma$ corresponding to Sierpiński space. As a topological space this is $\{\bot,\top\}$ with the topology generated by the single open $\{\top\}$. It can be formally defined as the locale corresponding to the free frame on a single generator (corresponding to $\{\top\}$). Locale maps $u\colon X\to\Sigma$ correspond to opens of $X$ by the universal property of the free frame. Spatially, this can be understood as taking the preimage of $\{\top\}$. From the verifiability perspective this allows us to interpret the elements of $\Sigma$ as corresponding to whether the verification process halts ($\top$), or runs forever ($\bot$).
Note that if $G$ is a set then $\Sigma^{G}$ represents the product $\Pi_{g\in G}\Sigma$. (We can alternatively describe it as an exponential object as defined in the following subsection, with $G$ viewed as a discrete locale.) The corresponding frame $\mathcal{O}(\Sigma^{G})$ is precisely $F(G)$, the free frame on $G$. So if $X$ is a locale with $\mathcal{O}X=\langle G\mid R\rangle$, we can dualise the coequaliser for $\langle G\mid R\rangle$ to obtain the locale equaliser
$X\hookrightarrow\Sigma^{G}\rightrightarrows\Sigma^{R}.$
@@
As mentioned before, compactness has a very nice interpretation in terms of universal quantification of verifiable properties. We will explore this topic further in Section˜4.
@@ 2.2. The Scott topology and exponentials
A subset $V\subseteq\mathcal{O}X$ is Scott-open if it is upward closed and if whenever a join $\bigvee D$ of a directed set $D$ is in $V$, then some $d\in D$ lies in $V$. Recall that a set $D$ is directed if every finite subset of $D$ has an upper bound in $D$. ∎
The Scott topology can actually be defined more generally on dcpos, which are often used to model data types in programming languages (as discussed in Section˜4.3).
@@
A poset which admits joins of all directed subsets is called a dcpo. We will write $\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}S$ for the join of a directed set $S$. A morphism of dcpos or a Scott-continuous function is a monotone map which preserves directed suprema. ∎
The order on a dcpo can be thought of as a definability order: the least element (if it exists) corresponds to a purely divergent computation, while maximal elements are completely defined. For more on this topic see [1, 2, 4, 11].
Definition 2.7.
The way-below relation $\ll$ on a dcpo $L$ is defined so that $a\ll b$ if and only if whenever $b\leq\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}D$ for a directed set $D$, there is some $d\in D$ with $a\leq d$. We say $L$ is continuous if for every $b\in L$ we have $b=\operatorname*{\bigvee{}^{\!\uparrow}_{\hskip-0.60275pt\hskip 2.2604pt}}\{a\in L\mid a\ll b\}$. (Note, in particular, that the set $\{a\in L\mid a\ll b\}$ is required to be directed).
Recall if $X,B$ are objects in category that the exponential object $B^{X}$ is equipped with a morphism $\mathrm{ev}\colon B^{X}\times X\to B$ such that for every $f\colon Y\times X\to B$ there is a unique $\overline{f}\colon Y\times X\to B^{X}$ such that the following diagram commutes.
$Y\times X$$^{X}\times XBB$$\overline{f}\times X$$\mathrm{ev}$$f$
The exponential is thought of as the object of functions from $X$ to $B$ and $\mathrm{ev}$ as the evaluation map. In the diagram above $\overline{f}$ can be viewed as the curried version of $f$. We say $B$ is exponentiable if the exponential $B^{X}$ exists for all objects $X$.
If $X$ is a locally compact locale, then $X$ is exponentiable (see [13]). In particular, the exponential locale $\Sigma^{X}$ exists and coincides with the Scott topology on $\mathcal{O}X\cong\mathrm{Hom}(X,\Sigma)$. Furthermore, $\Sigma^{X}$ is itself locally compact (for instance, see [28, Theorem 11.1]). However, if $X$ is not locally compact, the evaluation map $\mathrm{ev}\colon\mathcal{O}X\times X\to\Sigma$, which sends $(f,x)$ to $f(x)$, is not continuous when $\mathcal{O}X$ is given the Scott topology. In fact, in this case the exponential $\Sigma^{X}$ does not exist.
@@ 3. Machine space
We first tackle the problem of exponentials. We construct a ‘canonical weak exponential’ associated to a frame presentation and relate these to true exponentials when they exist. Along the way, we discuss the philosophical reasons behind why weak exponentials with base $\Sigma$ exist, but true exponentials might not.
3.1. Presentations and weak exponentials
For a general space $X$ the ‘space of opens’ $\Sigma^{X}$ need not exist. However, given a presentation $\mathcal{O}X=\langle G\mid R\rangle$ we will see that there is always a natural weak exponential given by $\Sigma^{\Sigma^{G}}$.
Recall that $\langle G\mid R\rangle$ is a quotient of the free frame on $G$, which in turn is the frame of opens of $\Sigma^{G}$. The frame quotient $\Sigma^{G}\twoheadrightarrow\langle G\mid R\rangle$ exhibits $X$ as a sublocale of $\Sigma^{G}$. The points of $\Sigma^{G}$ correspond to subsets of $G$ and we can then interpret this inclusion as sending points of $X$ to the set of generating opens in which they lie. The points in the image of this inclusion correspond to the frame homomorphisms $\mathcal{O}(\Sigma^{G})\to\mathcal{O}1$ which respect the relations. The additional points in $\Sigma^{G}$ may be thought of as ‘imaginary points’ in the same sense that the exponential spaces used in (the different approach of) [7, Chapter 10] are described as imaginary exponentials. Just as complex numbers may used to more easily prove results about real numbers, so will these imaginary points be useful for our study of compactness in Section˜4.
The space $\Sigma^{G}$ is always locally compact and so we can form $\Sigma^{\Sigma^{G}}$ — the space of opens of $\Sigma^{G}$. Of course, $\Sigma^{\Sigma^{G}}$ is a spatial locale whose points are Scott-opens of $\mathcal{O}X$ and whose topology is again given by the Scott topology. Alternatively, we can view $\Sigma^{\Sigma^{G}}$ as the double powerlocale on $G$ (see [28, 30]), the topology of which is generated by opens ${\boxtimes}U$ for each $U\subseteq G$ where an element $m=\bigvee_{i\in I}\bigwedge_{j\in J_{i}}g_{j}$ of $\Sigma^{\Sigma^{G}}$ lies in ${\boxtimes}U$ if and only if there exists some $J_{k}$ such that $g_{j}\in U$ for each $j\in J_{k}$.
We now show that $\Sigma^{\Sigma^{G}}$ can be made into a weak exponential. (Compare the construction of weak exponentials in $\mathrm{Top}$ given in [21].)